A firm is considering introducing a new product. It forecasts incremental an- nual gross profits (sales minus costs) from the product of $91,250, $106,750, and $114,250 for the next three years, respectively.

The project requires the purchase of a factory for $108,500, which would be straight-line depreciated over its 6-year tax life. It is estimated that the factory can be sold for $44,500 at the end of the three-year life of the project. Starting the project requires an initial net working capital investment of $21,000, with further annual increases of $5,750 during the life of the project. All net working capital is fully recoverable at cost when the project terminates. The firm pays tax at a marginal rate of 25% and the appropriate cost of capital for the project is 12.9% per annum.

The projected Free Cash Flow is closest to _________ at the end of Year 2 and closest to _________ at the end of Year 3.

a. $78,833; $122,708

b. $60,750; $151,562

c. $60,750; $131,396

d. $78,833; $169,646

e. $84,583; $175,396
Answer #1

Free Cash flow (FCFF)= EBIT* (1-T)+ D - changes in NWC- capex

Each year depreciation =108500/6= 18083

FCFF year 2= (106750- 18083)* 75% + 18083- 5750

= 78,833

Factory's book value after 3 years = Orginial cost- accumulated dep for 3 years=108500- 3*18083=54250

While we are selling at 44,500 i.e. at a loss of 9750 leading to tax shield of 9750*25%= 2438

FCFF year 3 = (114250-18083)*75%+ 18083+ Working capital recovery+ Disposal of asset + Tax saving because of disposal of asset at lower than book value

= (114250-18083)*75%+ 18083+ (21000+5750+5750)+44500+2438

= 169646